OpenStudy (anonymous):

WILL MEDAL The number of State Parks in California is 36 more than one third the number of State Parks in Florida. There are 85 State Parks in California. Which of the following equations would you use to find the number of State Parks in Florida? Note: f represents the number of state parks in Florida. 3(f − 36) = 85 1 over 3f + 36 = 85 Quantity f minus 36 all over 3 equals 85 Quantity f plus 36 all over 3 equals 85

OpenStudy (anonymous):

A :)

OpenStudy (anonymous):

@nicholls5645

OpenStudy (anonymous):

1/3f + 36 = 85 1/3f = 85 - 36 1/3f = 49 f = 49 * 3 f = 147